The difference of two positive numbers is 14. The larger is twice the smaller decreased by 5. Find the two numbers. (Enter solutions from smallest to largest.)

Answer by jorel1380(3719)

n-m=14
n=2m-5
n=m+14
m+14=2m-5
m=19
n=33
The two numbers are 19 and 33..